PSTricks

News – 2010

new macros and bugfixes for the basic package pstricks

April 23, 2011 Package author(s): Herbert Voß

Contents

I. pstricks – package 3

1. General 3 2. pstricks.sty 3

2.1. New optional argument . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3. pstricks.tex (2.20– 2011/04/23) 3

3.1. Macro \psLoop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 4. The PostScript header files 4

4.1. pstricks.pro . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 4.2. pst-algparser.pro . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

5. \psforeach and \psForeach 4

6. List of all optional arguments for pstricks 4 References 4

Part I.

pstricks – package

1. General

There exists a new document class pst-doc for writing PSTricks documentations, like this news document. It depends on the KOMA-Script document class scrartcl. pst-doc defines a lot of special macros to create a good index. Take one of the already existing package documentation and look into the source file. Then it will be easy to understand, how all these macros have to be used.

When running pdflatex the title page is created with boxes and inserted with the macro \AddToShipoutPicture from the package eso-pic. It inserts the background title page image pst-doc-pdf to use directly pdflatex. When running latex the title page is created with PSTricks macros.This allows to use the Perl script pst2pdf or the package pst-pdf or auto-pst-pdf or any other program/package which supports PostScript code in the document.

2. pstricks.sty

2.1. New optional argument

3. pstricks.tex (2.20– 2011/04/23)

3.1. Macro \psLoop

PSTricks already knows \psforeach and \psForeach for loops. The new macro \psLoop 2.17 allows a loop without defining a variable:

\psLoop{n}{argument}

However, the internal TEX counter psLoopIndex can be used for own purposes.

PSTricks PSTricks PSTricks PSTricks  ^{1 \psLoop {4}{ PSTricks }}

PSTricks PSTricks PSTricks
A  B  C  D

1 \tabular {| c| c| c| c|} 2 \psLoop {3}{ PSTricks &}\\\ hline 3 A & B & C & D\\\ hline 4 \endtabular 1 \begin {pspicture }[ showgrid ](3,3)

2 \psLoop {4}{ % 3 \psdots (\ the \psLoopIndex ,\ the \psLoopIndex ) } 4 \end {pspicture }

4. The PostScript header files

4.1. pstricks.pro

4.2. pst-algparser.pro

5. \psforeach and \psForeach

6. List of all optional arguments for pstricks

Key Type Default

References

[1] Michel Goosens, Frank Mittelbach, Sebastian Rahtz, Denis Roegel, and Herbert Voß. The L^{A}TEX Graphics Companion. Addison-Wesley Publishing Company, Reading, Mass., 2007.

[2] Laura E. Jackson and Herbert Voß. Die Plot-Funktionen von pst-plot. Die TEXnische Komödie, 2/02:27–34, June 2002.

[3] Nikolai G. Kollock. PostScript richtig eingesetzt: vom Konzept zum praktischen Einsatz. IWT, Vaterstetten, 1989.

[4] Herbert Voß. Die mathematischen Funktionen von Postscript. Die TEXnische Komödie, 1/02:40–47, March 2002.

[5] Herbert Voss. PSTricks Support for pdf. http://PSTricks.de/pdf/pdfoutput.

phtml, 2002.

[6] Herbert Voß. L^{A}TEX Referenz. DANTE – lehmanns media, Heidelberg/Hamburg, 2. edition, 2010.

[7] Herbert Voß. PSTricks – Grafik für TEX und L^{A}TEX. DANTE – Lehmanns Media, Heidelberg/Hamburg, 6. edition, 2010.

[8] Michael Wiedmann and Peter Karp. References for TEX and Friends. http://www. miwie.org/tex-refs/, 2003.

Index

\AddToShipoutPicture, 3 auto-pst-pdf, 3

Class pst-doc, 3 scrartcl, 3

Counter psLoopIndex, 3

eso-pic, 3

File pst-doc-pdf, 3

latex, 3

Macro \AddToShipoutPicture, 3 \psForeach, 3 \psforeach, 3 \psLoop, 3

Package auto-pst-pdf, 3 eso-pic, 3 pst-pdf, 3

pdflatex, 3 Program

latex, 3 pdflatex, 3 pst2pdf, 3

\psForeach, 3 \psforeach, 3 \psLoop, 3 psLoopIndex, 3 pst-doc, 3 pst-doc-pdf, 3 pst-pdf, 3 pst2pdf, 3

scrartcl, 3